OpenStudy (johndp123456):

Will fan and medal. In a survey of 65 students about favorite​ sports, the results include 42 who like​ tennis, 11 who like both tennis and​ football, and 4 who like neither of the sports. How many students like​ football? How many students like only​ tennis? How many students like only​ football?

OpenStudy (asc.bchs):

well if you add the numbers together, you get 57, which means you have 65 - 57 = 8 people who we don't know. based on the question, we can assume those 8 like only football. that means 8 + 42 = 50 students like football, 42 (provided) like tennis, and 8 like only football. im not 100% on this, i hope i at least helped :)

OpenStudy (johndp123456):

thank you

OpenStudy (asc.bchs):

np :) let me know if that is right; i wouldnt rely on my answer! :p

OpenStudy (johndp123456):

not right

OpenStudy (sshayer):



OpenStudy (johndp123456):

what is for what how can there be only 31 of there is 42

OpenStudy (sshayer):

31+11=42

OpenStudy (johndp123456):

oh

OpenStudy (johndp123456):

how is there only 11 that like football

OpenStudy (sshayer):

you got it or have any question?

OpenStudy (sshayer):

11 like both

OpenStudy (sshayer):

so only tennis=42-11=31

OpenStudy (johndp123456):

i mean sorry I got it mixed up but your answer is right

OpenStudy (sshayer):

no game =4